# Upper Mid-Band Channel Measurements and Characterization at 6.75 GHz FR1(C) and 16.95 GHz FR3 in an Indoor Factory Scenario

### v4

@@ Abstract
At higher mmWave frequencies, Schmieder et al. [13] observed unique propagation characteristics in industrial settings. Their measurements at 28 GHz and 3.7 GHz showed that LOS signals experienced approximately 3 dB less attenuation than predicted by FSPL up to 20 m, indicating better propagation due to constructive reflections and multipath enhancements. In NLOS conditions at 10 m distance, the 28 GHz signal experienced more severe attenuation (6 dB greater loss than FSPL), while the 3.7 GHz signal showed even higher attenuation (10 dB greater loss than FSPL), demonstrating that lower frequencies suffered more from obstruction in this environment.
@@ III-A Path Loss
After synthesizing the omnidirectional PDPs using the method described in [32, 21], in Fig. 3, we observe a lower PLE at 6.75 GHz, measuring 1.39 in LOS and 1.78 in NLOS, compared to 16.95 GHz, where the PLE values are 1.75 in LOS and 2.11 in NLOS. The low PLE at 6.75 GHz may be attributed to the structural characteristics of the MakerSpace, which includes plasterboard or wooden partitions with open sections at the top or bottom, facilitating easier penetration of lower-frequency waves. The relatively low PLE in NLOS at 6.75 GHz suggests a quasi-line-of-sight (QLOS) scenario, as defined in [33], where partial obstructions allow for some signal diffraction around obstacles, yielding a PLE less than 2, unlike fully obstructed NLOS conditions. Additionally, the absence of permanent walls contributes to the reception of scattered power from multiple directions at NLOS locations. The frequency comparison in Table II shows lower PLE at FR3 frequencies compared to sub-THz (e.g., Omni NLOS PLE of 1.78 at 6.75 GHz vs. 2.74 at 142 GHz) indicating MPC-rich propagation at FR3.
3GPP does not include a CI FSPL model with a 1 m reference distance option for path loss in InF environments; instead, it uses the Floating Intercept (FI) model with non-physical parameters that can vary over many orders of magnitude, as specified in Table 7.4.1-1: Path Loss Models in [12]. The FI model varies the y-intercept as a fitting parameter and does not use a physical reference of the free-space propagation as in the CI FSPL models with a 1 m reference distance [34]. Furthermore, 3GPP classifies the factory environment for TX heights above the surrounding clutter as sparse clutter-high height (SH) and dense clutter-high height (DH). Model parameters for both classifications are shown in Table II.
@@ III-B RMS Delay Spread
The RMS DS provides a measure of the temporal spread of MPCs in an environment. A power threshold of 25 dB below the peak or 5 dB above the noise floor is used to evaluate the RMS DS. Table III presents the RMS DS at 6.75 GHz and 16.95 GHz, along with the 3GPP results from Table 7.5-6 Part-3: Channel Model Parameters for InF in [12]. The mean RMS DS for the InF environment is evaluated by taking the of the RMS DS from the omni PDP at each TX-RX location assuming an underlying log-normal distribution for the RMS DS. The log operation on the data scales the distribution to a normal distribution allowing the mean () and standard deviation () of the normal distribution to be evaluated, for a direct comparison with the 3GPP log-normal models [12]. Finally, and are used to compute the expectation () to express the RMS DS in nanoseconds.
@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles of MPCs in azimuth and elevation planes. RMS AS is evaluated using a spatial lobe threshold of 10 dB below the strongest pointing direction, using the 3GPP method detailed in [12]. The AS is commonly evaluated using the 3GPP method [12] (circular standard deviation of MPCs) or Fleury’s method (second central moment of the PAS) [29]. A comprehensive comparison between 3GPP and Fleury’s method in computing AS is presented in [29]. Table IV summarizes the AS observed for arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.71∘ at 6.75 GHz and 14.32∘ at 16.95 GHz. Wider ASAs of 66.38∘ and 50.40∘ are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP uses a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [12]. Measured ASAs are observed to decrease with frequency in both LOS and NLOS at 6.75 GHz and 16.95 GHz. Further, we found ASDs to be 35.89∘ in LOS and 45.93∘ in NLOS at 6.75 GHz, while ASDs of 22.91∘ in LOS and 34.77∘ in NLOS were observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39∘ in LOS and 38.9∘ in NLOS from 0.5 GHz to 100 GHz.

### v5

@@ III-A Path Loss
After synthesizing the omni PDPs using the method described in [31, 20], in Fig. 3, we observe a lower PLE at 6.75 GHz, measuring 1.39 in LOS and 1.78 in NLOS, compared to 16.95 GHz, where the PLE values are 1.75 in LOS and 2.11 in NLOS. The low PLE at 6.75 GHz may be attributed to the structural characteristics of the MakerSpace, which includes plasterboard or wooden partitions with open sections at the top or bottom, facilitating easier penetration of lower-frequency waves. The relatively low PLE in NLOS at 6.75 GHz suggests a quasi-line-of-sight (QLOS) scenario, as defined in [32], where partial obstructions allow for some signal diffraction around obstacles, yielding a PLE less than 2, unlike fully obstructed NLOS conditions. Additionally, the absence of permanent walls contributes to the reception of scattered power from multiple directions at NLOS locations. The frequency comparison in Table II shows lower PLE at FR3 frequencies compared to sub-THz (e.g., Omni NLOS PLE of 1.78 at 6.75 GHz vs. 2.74 at 142 GHz) indicating MPC-rich propagation at FR3.
3GPP does not include a CI FSPL model with a 1 m reference distance option for path loss in InF environments; instead, it uses the Floating Intercept (FI) model with non-physical parameters that can vary over many orders of magnitude, as specified in Table 7.4.1-1: Path Loss Models in [12]. The FI model varies the y-intercept as a fitting parameter and does not use a physical reference of the free-space propagation [33]. Furthermore, 3GPP classifies the factory environment for TX heights above the surrounding clutter as sparse clutter-high height (SH) and dense clutter-high height (DH). Model parameters for both classifications are shown in Table II.
@@ III-B RMS Delay Spread
The RMS DS provides a measure of the temporal spread of MPCs in an environment. A power threshold of 25 dB below the peak or 5 dB above the noise floor is used to evaluate the RMS DS. Table III presents the RMS DS at 6.75 GHz and 16.95 GHz, along with the 3GPP results from Table 7.5-6 Part-3: Channel Model Parameters for InF in [12]. For a direct comparison with the log-normally distributed 3GPP RMS DS models, the RMS DS for the InF environment is evaluated by first taking the of the RMS DS from the omni PDP at each TX-RX location [12]. The operation converts the Omni RMS DS data into a normally distributed data, allowing the computation of arithmetic mean () and sample standard deviation () of the log-transformed RMS DS values and a direct comparison with 3GPP models [12]. In simpler words, one would calculate the arithmetic mean by taking the values in the “Omni DS” column of Table V for a measurement frequency and calculating , where OmniDSi is the Omni RMS DS at each TX-RX location and is the total number of measurement locations. Similarly, sample standard deviation is compute as , which provides an unbiased estimate of the standard deviation. The and are obtained for the normally distributed data and to express the distribution back into a linear nanosecond scale, the expectation is calculated using and ().
@@ III-C RMS Angular Spread
The RMS AS shows the spatial dispersion of arrival or departure angles of MPCs in azimuth and elevation planes. RMS AS is evaluated using a spatial lobe threshold of 10 dB below the strongest pointing direction, using the 3GPP method detailed in [12]. The AS is commonly evaluated using the 3GPP method [12] (circular standard deviation of MPCs) or Fleury’s method (second central moment of the PAS) [28]. A comprehensive comparison between the 3GPP model and conventional Fleury’s method for computing AS is presented in [28]. Similarly, by preserving the underlying log-normal distribution of the measured AS, we compare our AS measurements with the 3GPP model using the same approach as that applied to DS, as explained in Section III. B. Table IV summarizes the AS observed for arrival angles at the RX (ASA) and departure angles at the TX (ASD). In LOS, the ASA is observed as 23.71∘ at 6.75 GHz and 14.32∘ at 16.95 GHz. Wider ASAs of 66.38∘ and 50.40∘ are observed in NLOS at 6.75 GHz and 16.95 GHz, respectively. 3GPP uses a frequency-dependent LOS ASA with overestimated spread, along with a constant NLOS ASA, as outlined in Table 7.5-6 Part-3: Channel Model Parameters for InF and Table 7.5-11: ZSD and ZOD Offset Parameters for InF in [12]. Measured ASAs are observed to decrease with frequency in both LOS and NLOS at 6.75 GHz and 16.95 GHz. Further, we found ASDs to be 35.89∘ in LOS and 45.93∘ in NLOS at 6.75 GHz, while ASDs of 22.91∘ in LOS and 34.77∘ in NLOS were observed at 16.95 GHz. As shown in Fig. 4, the CDFs of empirical data align with these trends, while 3GPP predictions suggest a constant ASD of 39∘ in LOS and 38.9∘ in NLOS from 0.5 GHz to 100 GHz.